How much volume (in L) will 9,499 dg of a substance having a density of 1.57 g/mL occupy?


1
Expert's answer
2021-09-30T03:02:49-0400

Solution:

Convert dg to g:

10 dg = 1 g

Hence,

(9499 dg) × (1 g / 10 dg) = 949.9 g


Density = Mass / Volume

Volume = Mass / Density

Hence,

Volume = (949.9 g) / (1.57 g/mL) = 605.032 mL = 605 mL


Convert mL to L:

1000 mL = 1 L

Hence,

(605 mL) × (1L / 1000 mL) = 0.605 L


Answer: 0.605 L